A woman's age and her son's age add up to 45,5years ago the woman was 6times older than her son. How old was the woman when the son was born
Answer by mananth(16949)

woman's age = x
son's age = 45-x
5 years ago
woman's age = x-5
son's age = 45-x-5=40-x
x-5=6(40-x)
x-5=240-6x
7x=245
x=245/7
x=35
son's age = 10 now.
when he was born ahe was 25 years